are the two equations -6 + y = 2x and 2y - 4x =12 dependent
Answer by ewatrrr(24785)

-6 + y = 2x and 2y - 4x =12
...
-2(2x - y = -6)
4x - 2y = -12
0 = 0 Infinite Solutions |Same Line: y= 2x + 6
dependent? YES
....
A consistent system has at least one solution.
An inconsistent system has no solutions.
An independent system has exactly one solution.
A dependent system has infinitely many solutions.